OpenStudy (daryan):

Solve 2cosx + sin2x = 0 in the interval [0,2pi)?

OpenStudy (anonymous):

\[2 \cos x+2 \sin x ~\cos x=0 ,2 \cos x \left( 1+\sin x \right)=0\] cos x=0,find x sin x=-1,find x